# FedP2EFT: Federated Learning to Personalize PEFT for Multilingual LLMs

### v2

@@ 3.3 FedP2EFT: FL to Personalize PEFT
Overfitting often occurs when training an effective per-client PS in isolation due to limited client data. Following FedL2P (Lee et al. 2023), we mitigate this by federatedly learning a common PSG that generates client-wise PS. Concretely, we use a one hidden layer multilayer perceptron (MLP) with parameters $\phi$ that takes as input the client meta-data and outputs an estimated PS:
@@
Federated Training of FedP2EFT. Fig. 2 shows the entire FedP2EFT algorithm during federated training. For each federated round, each sampled participating client $i$ receives $\phi$ from the server and loads them into its MLP. They then 1 perform a forward pass of the local train dataset on their base model and a forward pass of the MLP with the resulting features as per Eq. (3.3). 2 The estimated $\bm{\hat{\lambda}}^{i}$ is plugged into our proposed BT-LoRA (Section 3.2) and 3 fine-tuning is performed as per Eq. (1) for $s$ steps (Stage 1). 4 The resulting $\hat{\bm{\lambda}}^{i,s}$ is used as an approximated ground-truth for regression learning of MLP to target the learned $\hat{\bm{\lambda}}^{i,s}$, where $\mathcal{L}_{1}$ is the L1 loss (Stage 2). Finally, 5 $\phi$ is sent back to the server for aggregation. As there is no single aggregation method that outperforms all others in every situation (Matsuda et al. 2022; Chen et al. 2022; Ye et al. 2024), we utilize FedAvg (McMahan et al. 2017). The aggregated $\phi$ is then sent to clients for the next round. To complement Figure 2, we provide Algorithm 1 in Appendix D.
@@ 4 Evaluation
We conduct experiments on multilingual scenarios, where clients with diverse high- and low-resource languages can collaboratively learn how to personalize a given base model to better cater to their language preferences. In all experiments, we divide clients in two pools, seen and unseen, where only the clients in the seen pool actively participate in federated training. We set the maximum number of communication rounds for training the PSG to $150$, randomly sampling $10\%$ of participating clients every round. We use Adam (Kingma and Ba 2015) as the default optimizer for all our experiments. Unless mentioned otherwise, we show results for resource budget $r=2$ where the total rank budget is $r\cdot L$ and leave results for other resource budgets $r=4,8,16$ in the Appendix. In the following subsections, we summarize the FL scenarios that we consider in our experiments, leaving comprehensive details in Appendix A.
@@ 4.1 Setup: Tasks, Models, and Datasets
Text Classification. We adopt the pretrained multilingual BERT (Devlin et al. 2018) (mBERT) for all text classification experiments. For datasets, we introduce additional data heterogeneity to the simulated FL setups, XNLI (Conneau et al. 2018) and MasakhaNEWS (Adelani et al. 2023), proposed in PE_FL (Zhao et al. 2024).
For our XNLI setup, we sample 2k instances for train and $500$ for test in each pool. In contrast to PE_FL, which had $15$ clients ($1$ client per language), we divide the data equally among $20$ clients per language. We then adopt the latent Dirichlet allocation (LDA) partition method (Hsu, Qi, and Brown 2019; Yurochkin et al. 2019), $y\sim Dir(\alpha)$, to simulate non-IID label shifts among these clients, with $\alpha=0.5$. Hence, there is a total of $600$ clients ($15$ languages $\cdot$ 2020 clients $\cdot$ 22 pools), consisting of both label and feature heterogeneity.
@@
Instruction-Tuning Generation. We use pretrained MobileLLaMA-1.4B (Chu et al. 2023) and Llama-3.2-3B (Dubey et al. 2024), which are representative of commonly supported model sizes on high-end edge devices (Mehta et al. 2024; Tan et al. 2024; Xu et al. 2024a), and run experiments with them on the Fed-Aya dataset (Ye et al. 2024). Fed-Aya, a real-world FL dataset naturally partitioned by annotator ID, comprises 38 clients, $8$ of whom we designate as our unseen pool. We split each client’s data $80\%/20\%$ for training and testing. Each client has up to $4$ languages and Fig. 3 shows the predominant language (client’s most frequent language) distribution in our setup.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Table 2: Mean±SD Accuracy of each language for unseen clients of our MasakhaNEWS setup ($r=2$, see Appendix Table 8 for other $r$ values). The pretrained model is trained using Standard FL with full fine-tuning and the resulting base model is personalized to each client given a baseline approach.
@@ 4.2 Setup: Complementary Approaches
Standard FL. We further train the pretrained model federatedly on the seen pool, either using existing PEFT methods or full fine-tuning (Ye et al. 2024; Sun et al. 2024),
Personalized FL (pFL). We adopt two recent pFL works: i) FedDPA-T (Yang et al. 2024b), which learns per-client personalized LoRA modules in addition to global LoRA modules, and ii) DEPT (SPEC) (Iacob et al. 2025), which learns per-client personalized token and positional embeddings while keeping the rest of the model shared. The base model hence differs for each client.
@@ 4.3 Setup: Baselines
LoRA PEFT. We deploy LoRA (Hu et al. 2022) on all linear layers of the model with a fixed rank $r$.
Non-FL Rank Selection. We compare with AdaLoRA (Zhang et al. 2023) and our proposed LoRA-variant of BayesTune (Kim and Hospedales 2023), BT-LoRA (Section 3.2), which optimizes $\bm{\lambda}$ separately for each client.
FL to Personalize. We compare with FedL2P (Lee et al. 2023) which trains a MLP federatedly to output per-client learning rates for each LoRA module.
@@ 4.4 Results on Text Classification
We evaluate our approach in a typical FL setup, where the pretrained model is first trained using Standard FL with full fine-tuning and the resulting base model is then personalized to each client. Tables 1 & 2 show the mean and standard deviation (SD) of the accuracy for each language in our MasakhaNEWS setup for seen and unseen pool respectively (similarly for XNLI in Appendix Tables 13 & 14).
The results in all four tables show that federated learning to personalize methods (FedL2P and FedP2EFT) outperform the other baselines in most cases. Non-FL rank selection approaches (AdaLoRA and BT-LoRA), on the other hand, tend to overfit and/or struggle to learn an optimal rank structure given the limited number of samples in each client. Comparing FedL2P and FedP2EFT, FedP2EFT largely surpass FedL2P with a few exceptions, indicating that learning to personalize LoRA rank structure is the better hyperparameter choice than personalizing learning rates; this finding is also aligned with recent LLM-based optimizer findings (Zhao et al. 2025), which shows that Adam’s performance is robust with respect to its learning rate.
FedP2EFT’s pFL Compatibility. Apart from Standard FL, we show that FedP2EFT can be plugged into existing pFL works that trains both a subset of the pretrained model and personalized layers for each client. Table 3 and Appendix Table 12 show that FedP2EFT outperforms baselines in almost all cases in our XNLI setup given a base model trained using FedDPA-T (Yang et al. 2024b) and DEPT(SPEC) (Iacob et al. 2025) respectively. In short, FedP2EFT can be integrated into a larger family of existing pFL approaches, listed in Section 2, to further improve personalization performance.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Table 3: Mean±SD Accuracy of each language across 3 different seeds for clients in the seen pool of our XNLI setup ($r=2$, see Appendix Table 9 for other $r$ values). The pretrained model is trained using FedDPA-T and the resulting base model is personalized to each client given a baseline approach. See Appendix Table 12 for results with DEPT (SPEC).
@@ 4.5 Results on Instruction-Tuning Generation
We evaluate our approach on the more challenging real-world multilingual benchmark, Fed-Aya. Tables 4 and 5 show the average METEOR (Banerjee and Lavie 2005)/ROUGE-L (Lin 2004) [^5] Due to limited space, we include ROUGE-1 in the Appendix. of each language given the off-the-shelf instruction finetuned Llama-3.2-3B (Llama-3.2-3B-Instruct) for seen and unseen clients respectively. Similarly, in Appendix Tables 15 and 16, we show the same tables given a pretrained MobileLLaMA-1.4B model trained using Standard FL with LoRA following the training recipe from FedLLM-Bench (Ye et al. 2024). These two models represent scenarios where the base model may or may not be trained using FL.
@@
ii) Predominant language dependence. None of the clients in the unseen pool have fr as a predominant language (Fig. 3). Because the number of predominant language samples heavily influences $\bm{\lambda}$, the performance for fr is worse. Instance-level $\bm{\lambda}$ generation, though a solution, is too costly, requiring one PSG forward pass per sample. This calls for novel efficient solutions that can better handle each client’s minority languages and is left as future work.
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lang. LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@ 4.7 Further Analysis
While the rank structure is the same across languages, the rank-wise scales (absolute values of $\bm{\lambda}$) differ. Following FedL2P, we visualize the difference in $\bm{\lambda}$ for different languages using the normalized mean distance, $d(j,k)$, between all clients pairs holding data for languages $j$ and $k$. Fig. 4 and Appendix Fig. 15 show these distances for XNLI and MasakhaNEWS setup respectively. Specifically, the value of each block in each figure is computed as follows: $\log(\frac{d(j,k)}{\sqrt{d(j,j)}\sqrt{d(k,k)}})$. Hence, the smaller the distance, the more similar $\bm{\lambda}$ is between languages. The results are aligned with our intuition that similar languages have similar $\bm{\lambda}$. For instance, the closest language to Urdu (ur) is Arabic (ar), both of which have the closest $\bm{\lambda}$ similarity (Fig. 4); likewise, for Tigrinya (tir) and Amharic (amh) in Appendix Fig. 15. We also observe that unrelated languages have similar $\bm{\lambda}$, e.g., Mandarin (zh) and Vietnamese (vi) share similar $\bm{\lambda}$ with the Indo-European languages (Fig. 4). This finding adds to existing evidence that leveraging dissimilar languages can sometimes benefit particular languages (Ye et al. 2024).
@@ A.1 Tasks, Datasets, and Data Partitioning
XNLI (Conneau et al. 2018). A natural language inference benchmark dataset for evaluating cross-lingual understanding covering 15 diverse languages including both high- and low-resources languages: English, French, Spanish, German, Greek, Bulgarian, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, Hindi, Swahili and Urdu. XNLI consists of premise-hypothesis pairs, labeled as entailment, contradiction, or neutral across different languages. We sample 2k instances from the XNLI train split and 500 instances from the test split for each pool. The data is then divided equally among 20 clients for each language using the latent Dirichlet allocation (LDA) partition with $\alpha=0.5$. Hence, the total number of clients is 600 (15 languages $\cdot$ 20 clients per language $\cdot$ 2 pools).
MasakhaNEWS (Adelani et al. 2023). A news topic classification benchmark designed to address the lack of resources for African languages. It covers 2 high-resource languages, English and French, and 14 low-resource languages, namely Amharic, Hausa, Igbo, Lingala, Luganda, Naija, Oromo, Rundi, chiShona, Somali, Kiswahili, Tigrinya, isiXhosa, and Yorùbá. Each sample contains a headline, the body text, and one of the 7 labels: business, entertainment, health, politics, religion, sports, or technology. We first combine all samples from the MasakhaNEWS train and validation split to form our train set, and use the MasakhaNEWS test split as our test set. We then split both train and test in each of the 16 languages by half for each pool. Following our XNLI setup, we adopt LDA $\alpha=0.5$ and divide each language’s data equally into 10 clients. Hence, the total number of clients is 320 (16 languages $\cdot$ 10 clients per language $\cdot$ 2 pools). Note that unlike XNLI, the number of samples for each language differs, hence there is quantity skew across clients.
Fed-Aya (Ye et al. 2024). A federated instruction tuning benchmark, based on Aya (Singh et al. 2024), where the data is annotated by contributors and partitioned by annotator ID. Following FedLLM-Bench (Ye et al. 2024), we focus on 6 high-resource languages, English, Spanish, French, Russian, Portuguese, Chinese, and 2 low-resource languages, Arabic and Telugu. Additionally, we filter out the other languages from the dataset. Out of 38 clients, we select 8 for our unseen pool, $\text{client\_ids}=[21,22,23,24,25,26,27,34]$ and the rest goes into our seen pool. Each client can have up to 4 languages where the number of data samples can range from a hundred to over a thousand samples per client.
@@ A.2 Models, Tokenizers, and Data Preprocessing
mBERT (Devlin et al. 2018). We use the pretrained multilingual BERT with its WordPiece tokenizer for all sequence classification experiments, namely all XNLI and MasakhaNEWS setups with various base models. For both datasets, we use a training batch size of 32 and pad input tokens on the right to a max token length of 128 and 256 respectively.
MobileLLaMA-1.4B (Chu et al. 2023). We train a base model with a pretrained MobileLLaMA-1.4B with Standard FL using LoRA in our Fed-Aya setup. We use the default LLaMA tokenizer which is a BPE model based on sentencepiece (Kudo and Richardson 2018) and adopt the UNK token as the PAD token. During training, we use an effective batch size of $16$ and pad right to the longest token in the batch up to a max token length of 1024. For evaluation, we use a batch size of 8, padding left instead, with greedy sampling up to a max new token length of 1024. We use the Alpaca template to format each prompt:
@@
Llama-3.2-3B (Dubey et al. 2024). We use the off-the-shelf Llama-3.2-3B-Instruct model as our base model and its default tokenizer which is a BPE model based on tiktoken [^7] https://github.com/openai/tiktoken. Training and evaluation hyperparameters are the same as the ones we use for MobileLLaMA. The only two differences are 1) we add a PAD token ‘<pad>’, and 2) we use the Llama 3 instruction template instead:

### v3

@@ 3.3 FedP2EFT: FL to Personalize PEFT
Overfitting often occurs when training an effective per-client PS in isolation due to limited client data. Following FedL2P (royson2023fedl2p), we mitigate this by federatedly learning a common PSG that generates client-wise PS. Concretely, we use a one hidden layer multilayer perceptron (MLP) with parameters $\phi$ that takes as input the client meta-data and outputs an estimated PS:
@@
Federated Training of FedP2EFT. Fig. 2 shows the entire FedP2EFT algorithm during federated training. For each federated round, each sampled participating client $i$ receives $\phi$ from the server and loads them into its MLP. They then 1 perform a forward pass of the local train dataset on their base model and a forward pass of the MLP with the resulting features as per Eq. (3.3). 2 The estimated $\bm{\hat{\lambda}}^{i}$ is plugged into our proposed BT-LoRA (Section 3.2) and 3 fine-tuning is performed as per Eq. (1) for $s$ steps (Stage 1). 4 The resulting $\hat{\bm{\lambda}}^{i,s}$ is used as an approximated ground-truth for regression learning of MLP to target the learned $\hat{\bm{\lambda}}^{i,s}$, where $\mathcal{L}_{1}$ is the L1 loss (Stage 2). Finally, 5 $\phi$ is sent back to the server for aggregation. As there is no single aggregation method that outperforms all others in every situation (matsuda2022empirical; chen2022pfl; fedllm-bench), we utilize FedAvg (fedavg). The aggregated $\phi$ is then sent to clients for the next round. To complement Figure 2, we provide Algorithm 1 in Appendix D.
@@ 4 Evaluation
We conduct experiments on multilingual scenarios, where clients with diverse high- and low-resource languages can collaboratively learn how to personalize a given base model to better cater to their language preferences. In all experiments, we divide clients in two pools, seen and unseen, where only the clients in the seen pool actively participate in federated training. We set the maximum number of communication rounds for training the PSG to $150$, randomly sampling $10\%$ of participating clients every round. We use Adam as the default optimizer for all our experiments. Unless mentioned otherwise, we show results for resource budget $r=2$ where the total rank budget is $r\cdot L$ and leave results for other resource budgets $r=4,8,16$ in the Appendix. In the following subsections, we summarize the FL scenarios that we consider in our experiments, leaving comprehensive details in Appendix A.
@@ 4.1 Setup: Tasks, Models, and Datasets
Text Classification. We adopt the pretrained multilingual BERT (BERT) (mBERT) for all text classification experiments. For datasets, we introduce additional data heterogeneity to the simulated FL setups, XNLI (XNLI) and MasakhaNEWS (MasakhaNEWS), proposed in PE_FL (zhao2023breaking).
For our XNLI setup, we sample 2k instances for train and $500$ for test in each pool. In contrast to PE_FL, which had $15$ clients ($1$ client per language), we divide the data equally among $20$ clients per language. We then adopt the latent Dirichlet allocation (LDA) partition method (hsu2019measuring), $y\sim Dir(\alpha)$, to simulate non-IID label shifts among these clients, with $\alpha=0.5$. Hence, there is a total of $600$ clients ($15$ languages $\cdot$ 2020 clients $\cdot$ 22 pools), consisting of both label and feature heterogeneity.
@@
Instruction-Tuning Generation. We use pretrained MobileLLaMA-1.4B (mobilellama) and Llama-3.2-3B (llama3), which are representative of commonly supported model sizes on high-end edge devices (2024_mobilequant) [^5] Refer to Appendix Sec E for experiments with Llama-3.1-8B., and run experiments with them on the Fed-Aya dataset (fedllm-bench). Fed-Aya, a real-world FL dataset naturally partitioned by annotator ID, comprises 38 clients, $8$ of whom we designate as our unseen pool. We split each client’s data $80\%/20\%$ for training and testing. Each client has up to $4$ languages and Fig. 3 shows the predominant language (client’s most frequent language) distribution in our setup.
Lan LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lan LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Table 2: Mean±SD Accuracy of each language for unseen clients of our MasakhaNEWS setup ($r=2$, see Appendix Table 8 for other $r$ values). The pretrained model is trained using Standard FL with full fine-tuning and the base model is personalized to each client given a baseline approach.
@@ 4.2 Setup: Complementary Approaches
Standard FL. We further train the pretrained model federatedly on the seen pool, either using existing PEFT methods or full fine-tuning (fedllm-bench; fedpeft),
Personalized FL (pFL). We adopt two recent pFL works: i) FedDPA-T (FedDPA), which learns per-client personalized LoRA modules in addition to global LoRA modules, and ii) DEPT (SPEC) (DEPT), which learns per-client personalized token and positional embeddings while keeping the rest of the model shared. The base model hence differs for each client.
@@ 4.3 Setup: Baselines
LoRA PEFT. We deploy LoRA (hu2021lora) on all linear layers of the model with a fixed rank $r$.
Non-FL Rank Selection. We compare with AdaLoRA (adalora) and our proposed LoRA-variant of BayesTune (kim2023bayestune), BT-LoRA (Section 3.2), which optimizes $\bm{\lambda}$ separately for each client.
FL to Personalize. We compare with FedL2P (royson2023fedl2p) which trains a MLP federatedly to output per-client learning rates for each LoRA module.
@@ 4.4 Results on Text Classification
We evaluate our approach in a typical FL setup, where the pretrained model is first trained using Standard FL with full fine-tuning and the resulting base model is then personalized to each client. Tables 1 & 2 show the mean and standard deviation (SD) of the accuracy for each language in our MasakhaNEWS setup for seen and unseen pool respectively (similarly for XNLI in Appendix Tables 15 & 16).
The results in all four tables show that federated learning to personalize methods (FedL2P and FedP2EFT) outperform the other baselines in most cases. Non-FL rank selection approaches (AdaLoRA and BT-LoRA), on the other hand, tend to overfit and/or struggle to learn an optimal rank structure given the limited number of samples in each client. Comparing FedL2P and FedP2EFT, FedP2EFT largely surpass FedL2P with a few exceptions, indicating that learning to personalize LoRA rank structure is the better hyperparameter choice than personalizing learning rates; this finding is also aligned with recent LLM-based optimizer findings (zhao2025deconstructing), which shows that Adam’s performance is robust with respect to its learning rate.
FedP2EFT’s pFL Compatibility. Apart from Standard FL, we show that FedP2EFT can be plugged into existing pFL works that trains both a subset of the pretrained model and personalized layers for each client. Table 3 and Appendix Table 14 show that FedP2EFT outperforms baselines in almost all cases in our XNLI setup given a base model trained using FedDPA-T (FedDPA) and DEPT(SPEC) (DEPT) respectively. In short, FedP2EFT can be integrated into a larger family of existing pFL approaches, listed in Section 2, to further improve personalization performance.
Lan LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Table 3: Mean±SD Accuracy of each language across 3 seeds for clients in the seen pool of our XNLI setup ($r=2$, see Appendix Table 9 for other $r$ values). The pretrained model is trained using FedDPA-T and the resulting base model is personalized to each client given a baseline approach. See Appendix Table 14 for results with DEPT (SPEC).
@@ 4.5 Results on Instruction-Tuning Generation
We evaluate our approach on the more challenging real-world multilingual benchmark, Fed-Aya. Tables 4 and 5 show the average METEOR (meteor)/ROUGE-L (ROUGE) [^6] Due to limited space, we include ROUGE-1 in the Appendix. of each language given the off-the-shelf instruction finetuned Llama-3.2-3B (Llama-3.2-3B-Instruct) for seen and unseen clients respectively. Similarly, in Appendix Tables 17 and 18, we show the same tables given a pretrained MobileLLaMA-1.4B model trained using Standard FL with LoRA following the training recipe from FedLLM-Bench (fedllm-bench). These two models represent scenarios where the base model may or may not be trained using FL.
@@
ii) Predominant language dependence. None of the clients in the unseen pool have fr as a predominant language (Fig. 3). Because the number of predominant language samples heavily influences $\bm{\lambda}$, the performance for fr is worse. This predominant language bias is quantified and future potential directions are discussed in Appendix Section E.
Lan LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@
Lan LoRA AdaLoRA BT-LoRA FedL2P FedP2EFT
@@ 4.7 Further Analysis
While the rank structure is the same across languages, the rank-wise scales (absolute values of $\bm{\lambda}$) differ. Following FedL2P, we visualize the difference in $\bm{\lambda}$ for different languages using the normalized mean distance, $d(j,k)$, between all clients pairs holding data for languages $j$ and $k$. Fig. 4 and Appendix Fig. 15 show these distances for XNLI and MasakhaNEWS setup respectively. Specifically, the value of each block in each figure is computed as follows: $\log(\frac{d(j,k)}{\sqrt{d(j,j)}\sqrt{d(k,k)}})$. Hence, the smaller the distance, the more similar $\bm{\lambda}$ is between languages. The results are aligned with our intuition that similar languages have similar $\bm{\lambda}$. For instance, the closest language to Urdu (ur) is Arabic (ar), both of which have the closest $\bm{\lambda}$ similarity (Fig. 4); likewise, for Tigrinya (tir) and Amharic (amh) in Appendix Fig. 15. We also observe that unrelated languages have similar $\bm{\lambda}$, e.g., Mandarin (zh) and Vietnamese (vi) share similar $\bm{\lambda}$ with the Indo-European languages (Fig. 4). This finding adds to existing evidence that leveraging dissimilar languages can sometimes benefit particular languages (fedllm-bench).
@@ A.1 Tasks, Datasets, and Data Partitioning
XNLI (XNLI). A natural language inference benchmark dataset for evaluating cross-lingual understanding covering 15 diverse languages including both high- and low-resources languages: English, French, Spanish, German, Greek, Bulgarian, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, Hindi, Swahili and Urdu. XNLI consists of premise-hypothesis pairs, labeled as entailment, contradiction, or neutral across different languages. We sample 2k instances from the XNLI train split and 500 instances from the test split for each pool. The data is then divided equally among 20 clients for each language using the latent Dirichlet allocation (LDA) partition with $\alpha=0.5$. Hence, the total number of clients is 600 (15 languages $\cdot$ 20 clients per language $\cdot$ 2 pools).
MasakhaNEWS (MasakhaNEWS). A news topic classification benchmark designed to address the lack of resources for African languages. It covers 2 high-resource languages, English and French, and 14 low-resource languages, namely Amharic, Hausa, Igbo, Lingala, Luganda, Naija, Oromo, Rundi, chiShona, Somali, Kiswahili, Tigrinya, isiXhosa, and Yorùbá. Each sample contains a headline, the body text, and one of the 7 labels: business, entertainment, health, politics, religion, sports, or technology. We first combine all samples from the MasakhaNEWS train and validation split to form our train set, and use the MasakhaNEWS test split as our test set. We then split both train and test in each of the 16 languages by half for each pool. Following our XNLI setup, we adopt LDA $\alpha=0.5$ and divide each language’s data equally into 10 clients. Hence, the total number of clients is 320 (16 languages $\cdot$ 10 clients per language $\cdot$ 2 pools). Note that unlike XNLI, the number of samples for each language differs, hence there is quantity skew across clients.
Fed-Aya (fedllm-bench). A federated instruction tuning benchmark, based on Aya (singh2024aya), where the data is annotated by contributors and partitioned by annotator ID. Following FedLLM-Bench (fedllm-bench), we focus on 6 high-resource languages, English, Spanish, French, Russian, Portuguese, Chinese, and 2 low-resource languages, Arabic and Telugu. Additionally, we filter out the other languages from the dataset. Out of 38 clients, we select 8 for our unseen pool, $\text{client\_ids}=[21,22,23,24,25,26,27,34]$ and the rest goes into our seen pool. Each client can have up to 4 languages where the number of data samples can range from a hundred to over a thousand samples per client.
@@ A.2 Models, Tokenizers, and Data Preprocessing
mBERT (BERT). We use the pretrained multilingual BERT with its WordPiece tokenizer for all sequence classification experiments, namely all XNLI and MasakhaNEWS setups with various base models. For both datasets, we use a training batch size of 32 and pad input tokens on the right to a max token length of 128 and 256 respectively.
MobileLLaMA-1.4B (mobilellama). We train a base model with a pretrained MobileLLaMA-1.4B with Standard FL using LoRA in our Fed-Aya setup. We use the default LLaMA tokenizer which is a BPE model based on sentencepiece (Kudo2018SentencePieceAS) and adopt the UNK token as the PAD token. During training, we use an effective batch size of $16$ and pad right to the longest token in the batch up to a max token length of 1024. For evaluation, we use a batch size of 8, padding left instead, with greedy sampling up to a max new token length of 1024. We use the Alpaca template to format each prompt:
@@
Llama-3.2-3B (llama3). We use the off-the-shelf Llama-3.2-3B-Instruct model as our base model and its default tokenizer which is a BPE model based on tiktoken [^8] https://github.com/openai/tiktoken. Training and evaluation hyperparameters are the same as the ones we use for MobileLLaMA. The only two differences are 1) we add a PAD token ‘<pad>’, and 2) we use the Llama 3 instruction template instead: